OpenStudy (anonymous):

find the general solution of the differential equation: y' + y(3/t) = t^2

OpenStudy (anonymous):

When I compute it, it becomes a real mess. So maybe you can help me?

OpenStudy (anonymous):

Are you having trouble with the homogeneous solution or the particular solution or...?

OpenStudy (anonymous):

I have problems with: finding K' if you see what I mean

OpenStudy (anonymous):

\(K'\) is likely a book specific thing. What else is it called?

OpenStudy (anonymous):

Actually I have no problems with finding y and y'. But then when I place what I got for y and y' in y' + y(3/t) = t^2 and try to find the function K(x) (this comes from the constant C) i get into trouble...

OpenStudy (anonymous):

What did you get for \(y(t)\)?

OpenStudy (anonymous):

(Kt)^-3

OpenStudy (anonymous):

K being e^C

OpenStudy (anonymous):

So that is the solution to the homogeneous equation: \[ y' + y(3/t) =0 \]Does that make sense?

OpenStudy (anonymous):

I'm not saying anything to help, just making sure we're on the same page here.

OpenStudy (anonymous):

yes that's right

OpenStudy (anonymous):

Now we want to find the particular solution. We need to make a guess at what \(y(t)\) would be. Notice that we know it is some sort of polynomial since \(t^2\) is a polynomial.

OpenStudy (anonymous):

So my guess would be \(At^n\), and we want to solve for \(n\) and \(A\). Does this guess make sense to you?

OpenStudy (anonymous):

I'm not sure but ok

OpenStudy (anonymous):

Yes, I know it sounds stupid, but we do have to actually guess at this stage.

OpenStudy (anonymous):

I dont understand how you come to that guess...?

OpenStudy (anonymous):

\[ y(t) = At^n \\ y'(t) = nAt^{n-1} \]

OpenStudy (anonymous):

Okay basically \(t^2\) is a polynomial, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

A polynomial plus a polynomial is a polynomial.

OpenStudy (anonymous):

The derivative of a polynomial is a polynomial.

OpenStudy (anonymous):

Oooookay I think your A is my K

OpenStudy (anonymous):

\(A\) is just some constant we need to solve for, so is \(n\).

OpenStudy (anonymous):

yes i understand

OpenStudy (anonymous):

\[ y(t) = At^n \\ y'(t) = nAt^{n-1} \]Plugging this in gets: \[ nAt^{n-1}+\frac{3At^n}{t}=t^2 \\ nAt^{n-1}+3At^{n-1}=t^2 \\ (nA+3A)t^{n-1}=t^2 \\ \]Okay so there are two neat equations here: \[ (nA+3A)t^{\color{red} {n-1}}=t^\color{red} {2} \implies n-1=2 \\ \color{red}{(nA+3A)}t^{n-1}=\color{red}{1}t^2 \implies nA+3A = 1 \\ \]

OpenStudy (anonymous):

We can see right away \(n=3\). Putting that into the second equaiton: \[ nA+3A = (3)A+3A = 1 \]

OpenStudy (anonymous):

Can you solve for \(A\)?

OpenStudy (anonymous):

1/6 sorry but I think I have to get a little bit sleep see you tomorrow

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

my brain is functioning to slowly

OpenStudy (anonymous):

It's correct.\[ \frac{1}{6}t^3 \]

OpenStudy (anonymous):

Full solution is \[ kt^{-3}+\frac{1}{6}t^3 \]

OpenStudy (anonymous):

But that isn't the answer, is it?

OpenStudy (anonymous):

It is a family of solutions because we need initial value like \(t(0)=1\) to solve for \(k\).

OpenStudy (anonymous):

that's right, i'll try to understand it tomorrow

OpenStudy (anonymous):

Whoops, I mean we need something like \(y(1) = 0\) or something.

OpenStudy (anonymous):

thanx a lot

OpenStudy (anonymous):

yes sure i didn't even notice

OpenStudy (anonymous):

Now that i've fully understood the concept I got the same result than you!!! Thanx for your help